How many grams of CuO would be required to produce 10.0 moles of nitrogen gas in the following chemical reaction?
2 NH₃ (g) + 3 CuO (s) → 3 Cu (s) + N₂ (g) + 3 H₂O (g)
1
Expert's answer
2019-10-09T07:35:35-0400

According to equation mole ratio "n(Cu):n(N_2)=3:1" , then "n(Cu) = 3\times n(N_2) = 3\times 10.0 = 30.0" mol


"m=M\times n"

"M(CuO) =79.55\frac{g}{mol}"

"m(CuO) = 30.0\times 79.55 = 2386.5 g = 2390 g"